OpenStudy (anonymous):

Find the amplitude y = 3 sin 1/4x

OpenStudy (johnweldon1993):

y = a * sin(bx - c) Where 'a' is the amplitude you have y = 3 *sin(1/4x) what is your amplitude...?

OpenStudy (anonymous):

3, I don't know

OpenStudy (johnweldon1993):

Yeah it's 3 :) That is the standard equation for a sin curve I wrote above y = a*sin(bx - c) 'a' is the amplitude of the function (how far vertically the graph reaches in both positive and negative direction) 'b' is the 'period' of the function (how long it takes the sin curve to go through 1 full cycle) and 'c' is called the phase shift (if the graph is moved left or right)

OpenStudy (anonymous):

oh, thanks:)

OpenStudy (johnweldon1993):

No problem...I hope THAT helped lol :)